# RoadBEV: Road Surface Reconstruction in Bird’s Eye View

### v2

@@ Abstract
Road surface conditions, especially geometry profiles, enormously affect driving performance of autonomous vehicles. Vision-based online road reconstruction promisingly captures road information in advance. Existing solutions like monocular depth estimation and stereo matching suffer from modest performance. The recent technique of Bird’s-Eye-View (BEV) perception provides immense potential to more reliable and accurate reconstruction. This paper uniformly proposes two simple yet effective models for road elevation reconstruction in BEV named RoadBEV-mono and RoadBEV-stereo, which estimate road elevation with monocular and stereo images, respectively. The former directly fits elevation values based on voxel features queried from image view, while the latter efficiently recognizes road elevation patterns based on BEV volume representing discrepancy between left and right voxel features. Insightful analyses reveal their consistence and difference with perspective view. Experiments on real-world dataset verify the models’ effectiveness and superiority. Elevation errors of RoadBEV-mono and RoadBEV-stereo achieve 1.83 cm and 0.50 cm, respectively. The estimation performance improves by 50% in BEV based on monocular image. Our models are promising for practical applications, providing valuable references for vision-based BEV perception in autonomous driving. The code is released at https://github.com/ztsrxh/RoadBEV
@@ I Introduction
In recent years, the rapid development of unmanned ground vehicles (UGVs) has posed higher requirements for on-board perception systems. Real-time understanding of driving environment and conditions is vital for accurate motion planning and control [1, 2, 3]. For vehicles, road is the only contacting media with the physical world. Road surface conditions determine many vehicle characteristics and driving performance [4]. Road unevenness like bumps and potholes (as shown in Fig. 1 (a)) exacerbates ride experience of passenger vehicles, which is intuitively perceptible. Real-time road surface condition perception, especially the geometry elevation, immensely contributes to planning and control systems in enhancing safety and ride comfort [5, 6, 7].
Figure 1: Our motivation. (a) Our reconstruction methods in BEV outperform these in image view for both monocular and stereo configurations. (b) For depth estimation in image view, the searching direction is biased from road elevation direction. Road profile features are sparse in depth view. The pothole is not clearly identifiable. (c) In BEV, profile vibrations are precisely captured such as the pothole, roadside step and even the rut. Road elevation feature in vertical direction is denser and easier to recognize.
Comparing with other perception tasks in UGVs like segmentation and detection, road surface reconstruction (RSR) is an emerging technique gaining more attention recently. Similar with existing perception pipelines, RSR generally utilizes on-vehicle LiDAR and camera sensors to retain road surface information. LiDAR straightforwardly scans road profiles and derives point clouds [8, 9]. Road elevation on vehicle trajectory can be directly extracted without complicated algorithms. Nevertheless, LiDAR sensors cost much, limiting their applications on economical mass-produced vehicles. Different from traffic objects with large size like vehicles and pedestrians, road unevenness generally have small amplitude and therefore, point cloud accuracy is essential. Motion compensation and filtering on real-time road scanning is necessary, which further requires high-accuracy localization at centimeter level.
Image-based RSR, a 3D vision task, is more promising than LiDAR in terms of accuracy and resolution. Road surface texture is also preserved, enabling more comprehensive road perception. Road elevation reconstruction based on vision is actually a depth estimation problem. For monocular camera, monocular depth estimation based on single image or multi-view stereo (MVS) based on sequence can be implemented to directly estimate depth [10]. For stereo camera, stereo matching regresses disparity map, which can be converted to depth [11, 12]. Given camera parameters, road point clouds in camera’s coordinate are recovered. Road structure and elevation information are finally obtained by primary post-processing pipelines. Supervised by ground-truth (GT) labels, high-accuracy and reliable RSR is achievable.
@@
Estimating road elevation from top-down view (i.e., BEV) is a natural idea, as elevation inherently describes vertical vibration. BEV is an efficient paradigm for representing multi-modal and multi-view data in an uniform coordinate [14, 15]. Recent SOTA performance on 3D object detection and segmentation tasks are achieved by BEV-based methods [16], which differs from the perspective view by introducing estimation heads on view-transformed image features. Fig. 1 shows our motivation. Instead of concentrating on global structure in image view, reconstruction in BEV directly recognize road features along vertical direction within a certain small range. Projected road features in BEV densely reflects structure and profile variation, contributing to effective and fine-grained searching. The influence of perspective effect is also suppressed, as road is uniformly represented on a plane vertical to view angle. Road reconstruction based on BEV features is expected to achieve higher performance.
In this paper, we reconstruct road surface in BEV to resolve the identified problems above. In particular, we focus on road geometry, i.e., elevation. To leverage monocular and stereo images and also, demonstrate the broad feasibility of BEV perception, we propose two sub-models named RoadBEV-mono and RoadBEV-stereo. Following the paradigm of BEV, voxels of interest covering potential road undulations are defined. The voxels query pixel features by 3D-2D projection. For RoadBEV-mono, elevation estimation head is introduced on reshaped voxel features. Structure of RoadBEV-stereo keeps consistent with stereo matching in image view. A 4D cost volume in BEV is built based on left and right voxel features, which is aggregated by 3D convolutions. Elevation regression is treated as classification on pre-defined bins for more efficient model learning. We validate the models on our previously released real-world dataset, showing their enormous superiority than the traditional monocular depth estimation an stereo matching methods.
@@ II Related Works
Road surface reconstruction by vision. Existing road reconstruction solutions are implemented in perspective view based on monocular or stereo images [17]. Early works recover road profile and detect anomaly by introducing prior geometry constraints [18, 19]. Based on known road surface disparity, road parallax or v-disparity map are built [20]. Affine transformation based on v-disparity is performed to localize irregular unevenness [21]. By introducing v-disparity road model and visual odometer, road elevation and drivable area are continuously extracted from stereo images. The geometry constraints above still rely on accurate disparity estimation. Road surface is sparsely reconstructed with structure from motion (SfM) based on sequence images in [22]. They assume local road surface is known and focus more on motion estimation with an adaptive Kalman Filter, which benefits global reconstruction in large-scale outdoor scenarios. The very recent research achieves large-scale monocular reconstruction via road mesh representations, recovering both geometry and texture [23, 24]. However, road elevation suffers from poor accuracy as they focus more on texture while geometry are supervised by sparse labels.
BEV representation.
BEV representation offers a coherent perspective for autonomous driving, facilitating accurate object localization and streamlined fusion of data from multiple sensors [25, 26]. This approach adeptly combines spatial and temporal data to enhance scene comprehension. Its applications span various real-world uses, including 3D object detection [27, 28, 29, 30, 31, 32], occupancy prediction [33, 34], motion planning [35], and online construction of HD maps [36]. Current works can be divided into two main categories based on view transformation: geometry-based and transformer-based. The transformer-based detectors, such as BEVFormer [30], follow a methodology where they first design a collection of BEV grid queries. These queries are then utilized to facilitate the view transformation through cross-attention with image features. The geometry-based approaches, such as Lift-splat-shoot (LSS) [37], involve lifting each image into a frustum of features based on depth [38, 28, 39] or height [31, 32] estimation and then splatting these frustums into a rasterized BEV grid. BEVDet [28] directly projects images into the BEV space for 3D object detection. Subsequent research endeavors introduce depth supervision from LiDAR sensors [29] or multi-view stereo techniques [40] to enhance depth estimation accuracy, leading to cutting-edge performance levels in this domain. There is a natural consistency in spatial distribution between the road surface and BEV grids, which makes the BEV paradigm naturally suitable for RSR tasks. Building on the provision of horizontal information in BEV space, we further introduce height estimation, characterized by its dense distribution and ease of prediction by the network, to achieve accurate RSR.
@@ III Dataset and Pre-processing
We utilize our previously released dataset, named Road Surface Reconstruction Dataset (RSRD) [41], as a benchmark to test model performance. This is a large-scale and high-accuracy dataset special for road surface reconstruction purpose, providing 2,800 pairs of high-resolution stereo images, dense point cloud labels, as well as motion pose information in the dense subset. Unlike existing vision datasets for autonomous driving perception, it focuses only on road surface and retains rich road textures. It covers diverse conditions of asphalt and concrete roads, including typical even and uneven cases like pothole and speed bump. We adopt the subset with half resolution (i.e., 960*540).For developing more reliable models and showcasing the significance of this application task, we extract representative samples with severer unevenness from the original dataset. The dataset utilized in this paper contains 1,210 training samples and 371 testing samples.
We first introduce the coordinate definitions shown in Fig. 2 (a). The $X_{c}\text{-}Y_{c}\text{-}Z_{c}$ is the original camera coordinate with a certain pitch angle w.r.t. the horizontal plane. The $X^{\prime}_{c}\text{-}Y^{\prime}_{c}\text{-}Z^{\prime}_{c}$ is an horizontal reference coordinate with $X^{\prime}_{c}\text{-}Z^{\prime}_{c}$ in the horizontal plane and $Y^{\prime}_{c}$ pointing the vertical direction (i.e., zero roll and pitch angles). The original and reference coordinates can transform to each other by utilizing the pose measured by IMU. Like depth estimation in image view where the camera plane indicates zero reference, road elevation also requires a reference base to describe profile properly. To facilitate algorithm development and the subsequent applications, we introduce another road coordinate $X_{r}\text{-}Y_{r}\text{-}Z_{r}$ vertically below the camera reference coordinate. The $X_{r}$ and $Y_{r}$ axes, which are parallel to $X^{\prime}_{c}$ and $Z^{\prime}_{c}$ respectively, represent the road lateral and longitudinal directions. The axis $Z_{r}$ now describes road profile, producing positive elevation values for roads above the reference plane; otherwise, negative. Based on our statistic analysis on the dataset, we set the referent height between $O^{\prime}_{c}$ and $O_{r}$ (i.e., the vertical distance between camera and road reference planes) as 1.10m. Although the exact camera-road distance changes due to vehicle suspension compression and elongation, the variation is small and the feature voxel defined below can cover the slight change.
@@
Figure 3: Examples of road image and GT elevation map.
@@ IV-A Feature Voxel and Elevation Regression
In the paradigm of BEV perception, 3D voxels are first defined to facilitate the transformation and feature projection from perspective view to 3D view. For the mainstream detection or segmentation tasks, feature voxels generally cover large range at hundred meters with big intervals in the three dimensions. Nevertheless, for this RSR task whose scale is much smaller, voxel interval should also be reduced to ensure high accuracy. For balancing accuracy and computation, we set the vertical voxel interval as 1.0 cm, which is smaller than the later and longitudinal intervals at 3.0 cm. This resolution covers the amplitude of slight road fluctuations like cracks and small rocks.
@@ IV-B RoadBEV-mono
Fig. 5 shows the architecture of RoadBEV-mono. The input RGB image undergoes a feature extraction backbone simplified from EfficientNet-B6 [42]. Similar with the common structures in detection, it comprises a feature pyramid with $\frac{1}{{2^{n}}}$ resolutions, where $n\in$1,2,3,4}. The multi-scale feature maps retain both low-level geometry and high-level semantic information. The feature pyramid is then interpolated to $\frac{1}{4}$ resolution, concatenated along the channel dimension, and finally fused as maps with $C$=64 channels. The corresponding indexing pixels of feature voxels are determined by projecting coordinates of voxel centers to the image plane. The voxels are then filled with the $C$-dimensional pixel features, resulting in voxel feature $\boldsymbol{F}_{vox}\in\mathbb{R}^{C\times N_{z}\times N_{y}\times N_{x}}$. As demonstrated in [43], to facilitate more efficient feature extraction for downstream tasks, the BEV feature $\boldsymbol{F}_{BEV}\in\mathbb{R}^{(C\cdot N_{z})\times N_{y}\times N_{x}}$ is derived by reshaping the vertical dimension of voxel feature. The BEV feature then undergoes 2D convolution with a reduced EfficientNet-B0, which is more cost-effective than the 3D.
@@
The mechanism analyzed above is actually the same with monocular depth estimation adopting search bins such as AdaBins [44]. Our RoadBEV differs from them as the search is directly along height direction instead of depth. It directly recognizes feature patterns of elevation variation, avoiding ineffective efforts in a biased direction. The BEV paradigm for RSR is thereby expected to be more effective than depth estimation.
@@ IV-C RoadBEV-stereo
Fig. 7 shows the architecture of RoadBEV-stereo. Similar with RoadBEV-mono and general stereo matching models, the left and right images first undergo feature extraction with shared weights. Settings about the CNN backbone are the same as RoadBEV-mono, i.e., EfficientNet-B6. The only difference is that the feature map resolution is interpolated to $\frac{1}{2}$ in the FPN. Explanations and analysis about resolution will be given below. The feature voxels are projected onto both the two image planes and query features from left and right perspectives. The left and right voxel features, $\boldsymbol{F}^{\prime}_{vox,l}$ and $\boldsymbol{F}^{\prime}_{vox,r}$, with the same shape $[C,N_{z},N_{y},N_{x}]$ are thereby obtained. In the architecture of stereo matching models, cost volume encoding similarities of the two feature maps is then built by correlation operation. We follow this paradigm and the voxel cost volume $\boldsymbol{V}_{discr}$ in BEV is derived as:
${\color[rgb]{0,0,0}\boldsymbol{V}_{discr}=\boldsymbol{F}^{\prime}_{vox,l}*\boldsymbol{F}^{\prime}_{vox,r}}$ (2)
@@
Similar with RoadBEV-mono, the 4D BEV volume can be reshaped as a 3D tensor and processed with 2D convolutions, which efficiently saves computation. Nevertheless, to demonstrate its consistence with stereo matching models, we still treat it as a 4D volume and aggregate by 3D convolutions. Specifically, we stack six 3D-conv layers and three hourglass modules with $\frac{1}{2}$ resolution reduction. The channel dimension of BEV volume is finally reduced to one. Since the number of vertical voxels $N_{z}$ may be different from the number of elevation classes $N_{c}$, the elevation proposal dimension is interpolated to have $N_{c}$ channels. The elevation feature map $\boldsymbol{F}^{\prime}_{ele}\in\mathbb{R}^{N_{c}\times N_{y}\times N_{x}}$ is thereby derived. The following output layers are the same with RoadBEV-mono.
Figure 7: Architecture of RoadBEV-stereo. The voxel defined under the left camera’s coordinate queries pixel features of the left and right feature maps. We construct cost volume in BEV by correlation between left and right voxel features. 3D convolutions then aggregate the 4D volume in BEV.

### v3

@@ Abstract
Road surface conditions, especially geometry profiles, enormously affect driving performance of autonomous vehicles. Vision-based online road reconstruction promisingly captures road information in advance. Existing solutions like monocular depth estimation and stereo matching suffer from modest performance. The recent technique of Bird’s-Eye-View (BEV) perception provides immense potential to more reliable and accurate reconstruction. This paper uniformly proposes two simple yet effective models for road elevation reconstruction in BEV named RoadBEV-mono and RoadBEV-stereo, which estimate road elevation with monocular and stereo images, respectively. The former directly fits elevation values based on voxel features queried from image view, while the latter efficiently recognizes road elevation patterns based on BEV volume representing correlation between left and right voxel features. Insightful analyses reveal their consistence and difference with the perspective view. Experiments on real-world dataset verify the models’ effectiveness and superiority. Elevation errors of RoadBEV-mono and RoadBEV-stereo achieve 1.83 cm and 0.50 cm, respectively. Our models are promising for practical road preview, providing essential information for promoting safety and comfort of autonomous vehicles. The code is released at https://github.com/ztsrxh/RoadBEV
Index Terms:
Bird’s eye view, 3D reconstruction, road surface condition, road preview, autonomous driving.
@@ I Introduction
In recent years, the rapid development of unmanned ground vehicles (UGVs) has posed higher requirements for on-board perception systems. Real-time understanding of driving environment and condition is vital for accurate motion planning and control [1, 2, 3]. For vehicles, road is the only contacting media with the physical world. Road surface conditions determine many vehicle characteristics and driving performance [4]. Road unevenness like bumps and potholes (as shown in Fig. 1 (a)) exacerbates ride experience of passenger vehicles, which is intuitively perceptible [5, 6]. Real-time road surface condition perception, especially the geometry elevation, immensely contributes to planning and control systems in enhancing safety and ride comfort [7, 8, 9].
Figure 1: Our motivation. (a) Our reconstruction methods in BEV outperform these in image view for both monocular and stereo configurations. (b) For depth estimation in image view, the searching direction is biased from road elevation direction. Road profile features are sparse in depth view. The pothole is not clearly identifiable. (c) In BEV, profile vibrations are precisely captured such as the pothole, roadside step and even the rut. Road elevation feature in vertical direction is denser and easier to be recognized.
Comparing with other perception tasks in UGVs like segmentation and detection, road surface reconstruction (RSR) is an emerging technique gaining more attention recently. Similar with existing perception pipelines, RSR generally utilizes on-vehicle LiDAR and camera sensors to retain road surface information. LiDAR straightforwardly scans road profile and derives point clouds [10, 11, 12]. Although road elevation on tire trajectories can be directly extracted without complicated post-processing, LiDAR sensors cost much, limiting their applications on economical vehicles.
Image-based RSR, a 3D vision task, is more promising than LiDAR in terms of accuracy and resolution. Road surface texture is also preserved, enabling more comprehensive road perception. Road geometry reconstruction based on vision is actually a depth estimation task. For monocular camera, monocular depth estimation based on single image or multi-view stereo (MVS) based on image sequence can be implemented to directly estimate depth [13]. For stereo camera, stereo matching regresses disparity map, which can be converted to depth [14, 15, 16]. Given camera parameters, road point clouds in camera’s coordinate are recovered. Road structure and elevation information are finally obtained by primary post-processing pipelines. Supervised by ground-truth (GT) labels, high-accuracy and reliable RSR is achievable.
@@
Estimating road elevation from top-down view (i.e., BEV) is a natural idea, as elevation inherently describes vertical vibration. BEV is an efficient paradigm for representing multi-modal and multi-view data in an uniform coordinate [18]. Recent SOTA performance on 3D object detection and segmentation tasks are achieved by BEV-based methods [19, 20, 21], which differs from the perspective view by introducing estimation heads on view-transformed image features. Fig. 1 shows our motivation. Instead of concentrating on global structure in image view, reconstruction in BEV directly recognize road features along vertical direction within a certain small range. Projected road features in BEV densely reflects structure and profile variation, contributing to effective and fine-grained searching. The influence of perspective effect is also suppressed, as road is uniformly represented on a plane vertical to view angle. Road reconstruction based on BEV features is expected to achieve higher performance.
In this paper, we reconstruct road surface in BEV to resolve the identified problems above. In particular, we focus on road geometry, i.e., elevation. To leverage monocular and stereo images and also, demonstrate the broad feasibility of BEV perception, we propose two sub-models named RoadBEV-mono and RoadBEV-stereo. Following the paradigm of BEV, voxels of interest covering potential road undulations are defined. The voxels query pixel features by 3D-2D projection. For RoadBEV-mono, elevation estimation head is introduced on reshaped voxel features. Structure of RoadBEV-stereo keeps consistent with the stereo matching models in image view. A 4D cost volume in BEV is built based on the left and right voxel features, which is further aggregated by 3D convolutions. Elevation regression is treated as classification on pre-defined bins for more efficient model learning. We validate the models on our previously released real-world dataset, showing their enormous superiority than the traditional monocular depth estimation and stereo matching methods.
@@ II Related Works
Road surface reconstruction by vision. Existing road reconstruction solutions are implemented in perspective view based on monocular or stereo images [22]. Early works recover road profile and detect anomaly by introducing prior geometry constraints [23, 24]. Based on known road surface disparity, road parallax or v-disparity map are built [25]. Affine transformation based on v-disparity is performed to localize irregular unevenness [26]. By introducing v-disparity road model and visual odometer, road elevation and drivable area are continuously extracted from stereo images. The geometry constraints above still rely on accurate disparity estimation. Road surface is sparsely reconstructed with structure from motion (SfM) based on sequence images in [27]. They focus more on motion estimation with an adaptive Kalman Filter, which benefits continuous and global reconstruction in large-scale outdoor scenarios. The very recent research achieves large-scale monocular reconstruction via road mesh representations, recovering both geometry and texture [28, 29]. However, road elevation suffers from poor accuracy as they focus more on texture while geometry are supervised by sparse labels.
BEV representation. BEV representation offers a coherent perspective for autonomous driving, facilitating accurate object localization and streamlined fusion of data from multiple sensors [30, 31, 32, 33]. This approach adeptly combines spatial and temporal data to enhance scene comprehension. Its applications span various real-world uses, including 3D object detection [34, 35, 36, 37, 38], occupancy prediction [39], motion planning [40], and online construction of HD maps [41]. Current works can be divided into two main categories based on view transformation: geometry-based and transformer-based. The transformer-based detectors, such as BEVFormer [42], follow a methodology where they first design a collection of BEV grid queries. These queries are then utilized to facilitate the view transformation through cross-attention with image features. The geometry-based approaches, such as Lift-splat-shoot (LSS) [43], involve lifting each image into a frustum of features based on depth [44] or height [36, 45] estimation and then splatting these frustums into a rasterized BEV grid. BEVDet [46] directly projects images into the BEV space for 3D object detection. Subsequent research endeavors introduce depth supervision from LiDAR sensors [47] or multi-view stereo techniques [48] to enhance depth estimation accuracy, leading to cutting-edge performance levels in this domain. There is a natural consistency in spatial distribution between the road surface and BEV grids, which makes the BEV paradigm naturally suitable for RSR tasks. Building on the provision of horizontal information in BEV space, we further introduce height estimation, characterized by its dense distribution and ease of prediction by the network, to achieve accurate RSR.
@@ III Dataset and Pre-processing
We utilize our previously released dataset, Road Surface Reconstruction Dataset (RSRD) [49], as a benchmark to test model performance. This is a large-scale and high-accuracy dataset special for road surface reconstruction purpose, providing 2,800 pairs of high-resolution stereo images, dense point cloud labels, as well as motion pose information in the dense subset. Unlike existing vision datasets for autonomous driving perception, it focuses only on road surface and retains rich road textures. It covers diverse conditions of asphalt and concrete roads, including typical even and uneven cases like pothole and speed bump. We adopt the subset with half resolution (i.e., 960*540). For developing more reliable models and showcasing the significance of this application task, we extract representative samples with severer unevenness from the original dataset. The dataset utilized in this paper contains 1,210 training samples and 371 testing samples.
We first introduce the coordinate definitions shown in Fig. 2 (a). The $X_{c}\text{-}Y_{c}\text{-}Z_{c}$ is the original camera coordinate with a certain pitch angle w.r.t. the horizontal plane. The $X^{\prime}_{c}\text{-}Y^{\prime}_{c}\text{-}Z^{\prime}_{c}$ is an horizontal reference coordinate with $X^{\prime}_{c}\text{-}Z^{\prime}_{c}$ in the horizontal plane and $Y^{\prime}_{c}$ pointing the vertical direction (i.e., zero roll and pitch angles). The original and reference coordinates can transform to each other with the pose measured by IMU. Like depth estimation in image view where the camera plane indicates zero reference, road elevation also requires a reference base to describe profile properly. To facilitate algorithm development and the subsequent applications, we introduce another road coordinate $X_{r}\text{-}Y_{r}\text{-}Z_{r}$ vertically below the camera reference coordinate. The $X_{r}$ and $Y_{r}$ axes, which are parallel to $X^{\prime}_{c}$ and $Z^{\prime}_{c}$ respectively, represent the road lateral and longitudinal directions. The axis $Z_{r}$ now describes road profile, producing positive elevation values for roads above the reference plane; otherwise, negative. Based on our statistic analysis on the dataset, we set the referent height between $O^{\prime}_{c}$ and $O_{r}$ (i.e., the vertical distance between camera and road reference planes) as 1.10 m. Although the exact camera-road distance changes due to vehicle suspension compression and elongation, the variation is small and the feature voxel defined below can cover the slight change.
@@
Figure 3: Examples of road image and GT elevation map. The unit of colorbar is cm.
@@ IV-A Feature Voxel and Elevation Regression
In the paradigm of BEV and occupancy perception, 3D voxels are first defined to facilitate the transformation and feature projection from perspective view to 3D view. For the mainstream detection or segmentation tasks, feature voxels generally cover large range at hundred meters with big intervals in the three dimensions. Nevertheless, for this RSR task whose scale is much smaller, voxel interval should also be reduced to ensure high accuracy. Balancing accuracy and computation, we set the vertical voxel interval as 1.0 cm, which is smaller than the lateral and longitudinal intervals of 3.0 cm. This resolution covers the amplitude of slight road fluctuations like cracks and small rocks.
@@ IV-B RoadBEV-mono
Fig. 5 shows the architecture of RoadBEV-mono. The input RGB image undergoes a feature extraction backbone simplified from EfficientNet-B6 [50]. Similar with the common structures in detection, it comprises a feature pyramid with $\frac{1}{{2^{n}}}$ resolutions, where $n\in$1,2,3,4}. The multi-scale feature maps retain both low-level geometry and high-level semantic information. The feature pyramid is then interpolated to $\frac{1}{4}$ resolution, concatenated along the channel dimension, and finally fused as maps with $C$=64 channels. The corresponding indexing pixels of feature voxels are determined by projecting coordinates of voxel centers to the image plane. The voxels are then filled with the $C$-dimensional pixel features, resulting in voxel feature $\boldsymbol{F}_{vox}\in\mathbb{R}^{C\times N_{z}\times N_{y}\times N_{x}}$. As demonstrated in [51], to facilitate more efficient feature extraction for downstream tasks, the BEV feature $\boldsymbol{F}_{BEV}\in\mathbb{R}^{(C\cdot N_{z})\times N_{y}\times N_{x}}$ is derived by reshaping the vertical dimension of voxel feature. The $\boldsymbol{F}_{BEV}$ then undergoes 2D convolution with a reduced EfficientNet-B0, which is more cost-effective than the 3D.
@@
The mechanism analyzed above is actually the same with monocular depth estimation adopting search bins such as AdaBins [52]. Our RoadBEV differs from them as the search is directly along height direction instead of depth. It directly recognizes feature patterns of elevation variation, avoiding ineffective efforts in a biased direction. The BEV paradigm for RSR is thereby expected to be more effective than depth.
@@ IV-C RoadBEV-stereo
Fig. 7 shows the architecture of RoadBEV-stereo. Similar with RoadBEV-mono and general stereo matching models, the left and right images first undergo feature extraction with shared weights. Settings about the CNN backbone are the same as RoadBEV-mono, i.e., EfficientNet-B6. The only difference is that the feature map resolution is interpolated to $\frac{1}{2}$ in the FPN. Explanations and analysis about this resolution setting will be given below. The feature voxels are projected onto both the two image planes and query features from left and right perspectives. The left and right voxel features, $\boldsymbol{F}^{\prime}_{vox,l}$ and $\boldsymbol{F}^{\prime}_{vox,r}$, with the same shape $[C,N_{z},N_{y},N_{x}]$ are thereby obtained. In the architecture of stereo matching models, cost volume encoding similarities of the two feature maps is then built by correlation operation. We follow this paradigm and the voxel cost volume $\boldsymbol{V}_{discr}$ in BEV is derived as:
$\boldsymbol{V}_{discr}=\boldsymbol{F}^{\prime}_{vox,l}*\boldsymbol{F}^{\prime}_{vox,r}$ (2)
@@
Similar with RoadBEV-mono, the 4D BEV volume can be reshaped as a 3D tensor and processed with 2D convolutions, which efficiently saves computation. Nevertheless, to demonstrate its consistence with stereo matching models, we keep it as a 4D volume and aggregate by 3D convolutions. In this way, the estimation head is actually an occupancy network, which predicts the possibility of being occupied by road surface. Specifically, we stack six 3D-conv layers and three hourglass modules with $\frac{1}{2}$ resolution reduction. The channel dimension of BEV volume is finally reduced to one. Since the number of vertical voxels $N_{z}$ may be different from the number of elevation classes $N_{c}$, the elevation proposal dimension is interpolated to have $N_{c}$ channels. The elevation feature map $\boldsymbol{F}^{\prime}_{ele}\in\mathbb{R}^{N_{c}\times N_{y}\times N_{x}}$ is thereby derived. The following output layers are the same with RoadBEV-mono. Different from general occupancy networks for scene perception where possibility is given for every voxel, the possibility here is normalized along the vertical dimension, indicating occupancy among elevation candidates at a horizontal grid.
Figure 7: Architecture of RoadBEV-stereo. The voxel defined under the left camera’s coordinate queries pixel features of the left and right feature maps. We construct cost volume in BEV by correlation between left and right voxel features. 3D convolutions then aggregate the 4D volume, predicting the occupancy possibility among vertical voxels.